OpenStudy (anonymous):

On a distant planet, Ralph ran into a creature with multiple arms and multiple legs. Each arm has 6 fingers and each leg has 9 toes. On any one of these creatures, the total number of fingers and toes, D, can be described with the equation D = 6x + 9y. In this equation, what does the variable x represent? A. the number of arms the creature has B. the number of legs the creature has C. the number of fingers the creature has D. the number of toes the creature has Please Help

OpenStudy (the_fizicx99):

A because it has 6 toes, suppose it has 3 arms its 6*3. D is both the fingers and toes together.

OpenStudy (anonymous):

9 toes 6 fingers

OpenStudy (the_fizicx99):

Yes, D = # of toes and fingers combined. 6x represent the fingers each arm have. Like i said before, if it has 4 arms, he'll have 24 fingers +9y.. y = number of toes per leg

OpenStudy (anonymous):

ok so its D? or A?

OpenStudy (the_fizicx99):

A! D is from the equation D = 6x + 9y! lol

OpenStudy (anonymous):

sorry lol thx!

OpenStudy (the_fizicx99):

Basically, suppose he has 2 arms and legs. It would be D = 6(2) + 9(2) D = 12 + 18 D = 30 so TOTAL amount would be 30